A 1800-Hz note is played with a 1350-Hz note. 1800 Hz is the _____ harmonic of the beat frequency.
Answer #1

The beat frequency is, fbear = f2 - fi = 1800-1350 = 450 Hz

the harmonic is

n=1800/450

=4th harmonic frequency

The beat frequency is, fbear = f2 - fi = 1800-1350 = 450 Hz